Serial numbers for a product are to be made using 4 letters followed by
4 numbers. If the letters are to be taken from the first 6 letters of the alphabet with no repeats and the numbers are taken from the digits 0−5 with no repeats, how many serial numbers can be generated So that the numbers are together ?

Answer by ikleyn(52786)

.

(6*5*4*3) * (6*5*4*3) = 360 * 360 = 129600.         ANSWER

First multiplier  (6*5*4*3)  is for  4-letters block made of  6  different non-repeating letters.

Second multiplier  (6*5*4*3)  is for  4-digits block made of  6 different non-repeating digits.